Circular No. 6746

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

CATACLYSMIC VARIABLE IN PEGASUS Y.-l. Qiu, Q.-y. Qiao, and J.-y. Hu, Beijing Astronomical Observatory

(BAO); and A. Esamdin, Urmqi Astronomical Station, report their discovery in the course of the BAO supernova survey of a new cataclysmic variable star at α = 21^{h}12^{m}29^{s}.54, δ = +12^{o}32′06′′.5 (equinox 2000.0). R magnitudes obtained with the 0.60-m telescope at Xinglong Station: Sept. 3 UT, [19.0; 18, [18.0; 20.52, 16.0; 21.49, 17.3. The POSS-I plate shows a very faint object at this postion, but nothing is present in the USNO A1.0 catalogue. A spectrogram of the variable was obtained with the 2.16-m telescope at Xinglong on Sept. 21.58. The spectrum shows typical features of cataclysmic variables during eruption, with a blue continuum superimposed by weak Hα emission (equivalent width 1.12 nm), Hβ emission in the core of the absorption, and weak He I lines at 587.6 nm.

XTE J1755–324 J. Paul, A. Goldwurm, E. Jourdain, J. P. Roques, R. Sunyaev, E. Chu-

razov, S. Trudolyubov, and M. Revnivtsev, on behalf of the SIGMA/Granat team (Service d’Astrophysique du CEA, Saclay; Centre d’Etude Spatiale des Rayonnements, Toulouse; Space Research Institute, Moscow), communicate: “During the first observation of a new campaign dedicated to the galactic center, the SIGMA/Granat telescope detected a hard x-ray source at α = 17^{h}52^{m}, δ = −32^{o}.4 (equinox 1950.0; uncertainty ± 5′). The position is consistent with the newly-discovered soft source XTE J1755–324 (IAUC 6710). The observation started on Sept. 16 and lasted ∼ 40 hr. Preliminary estimates of the source flux give 75 ± 17 mCrab in the range 40–75 keV and 34 ± 20 mCrab over 75–150 keV. This is the first evidence of a hard component in the source spectrum extending above 40 keV, and — combined with a soft blackbody component observed by XTE — strongly suggests a blackhole binary.”

COMET 43P/WOLF-HARRINGTON Total visual magnitude estimates: Sept. 1.10 UT, 12.6 (K. Hornoch,

Lelekovice, Czech Republic, 0.35-m reflector); 1.11, 13.2 (R. J. Bouma, Groningen, The Netherlands, 0.25-m reflector); 2.06, 12.5 (M. Plsek, Lelekovice, Czech Republic, 0.35-m reflector); 11.11, 13.0 (Bouma); 12.09, 12.6 (M. Meyer, Frauenstein, Germany, 0.25-m reflector).

1997 September 22  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green